another oatrocal asin cooke? (Round yor anww to 4 decimal places Points pensible: t This is attept t of 1 is atemye of
Answer #1

Total number of cookies in the bag = 9+5+4+7 = 25

Probability of picking oatmeal raisin cookie for the first time = 7/25

After eating one, there will be 24 cookies remaining in the bag out of which 6 are oatmeal raisin cookies.

So, probability of picking one more oatmeal raisin cookie = 6/24

So, total probability = (7/25)*(6/24) = 7/100 = 0.07